# Graph Neural Preconditioners for Iterative Solutions of Sparse Linear Systems

### v1

@@ Abstract
Preconditioning is at the heart of iterative solutions of large, sparse linear systems of equations in scientific disciplines. Several algebraic approaches, which access no information beyond the matrix itself, are widely studied and used, but ill-conditioned matrices remain very challenging. We take a machine learning approach and propose using graph neural networks as a general-purpose preconditioner. They show attractive performance for ill-conditioned problems, in part because they better approximate the matrix inverse from appropriately generated training data. Empirical evaluation on over 800 matrices suggests that the construction time of these graph neural preconditioners (GNPs) is more predictable than other widely used ones, such as ILU and AMG, while the execution time is faster than using a Krylov method as the preconditioner, such as in inner-outer GMRES. GNPs have a strong potential for solving large-scale, challenging algebraic problems arising from not only partial differential equations, but also economics, statistics, graph, and optimization, to name a few.
@@ 1 Introduction
Iterative methods are commonly used to solve large, sparse linear systems of equations $\mathbf{A}\mathbf{x}=\mathbf{b}$. These methods typically build a Krylov subspace, onto which the original system is projected, such that an approximate solution within the subspace is extracted. For example, GMRES [30], one of the most popularly used methods in practice, builds an orthonormal basis $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}$ of the $m$-dimensional Krylov subspace by using the Arnoldi process and defines the approximate solution $\mathbf{x}_{m}\in\mathbf{x}_{0}+\operatorname{span}(\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\})$, such that the residual norm $\|\mathbf{b}-\mathbf{A}\mathbf{x}_{m}\|_{2}$ is minimized.
@@
The effectiveness of Krylov methods heavily depends on the conditioning of the matrix $\mathbf{A}$. Hence, designing a good preconditioner is critical in practice. In some cases (e.g., solving partial differential equations, PDEs), the problem structure provides additional information that aids the development of a preconditioner applicable to this problem or a class of similar problems. For example, multigrid preconditioners are particularly effective for Poisson-like problems with a mesh. In other cases, little information is known beyond the matrix itself. An example is the SuiteSparse matrix collection [8], which contains thousands of sparse matrices for benchmarking numerical linear algebra algorithms. In these cases, a general-purpose (also called “algebraic”) preconditioner is desirable; examples include ILU [28], approximate inverse [7], and algebraic multigrid (AMG) [27]. However, it remains very challenging to design one that performs well universally.
@@
In this work, we propose to use neural networks as a general-purpose preconditioner. Specifically, we consider preconditioned Krylov methods for solving problems in the form $\mathbf{A}\mathbf{M}\mathbf{u}=\mathbf{b}$, where $\mathbf{M}\approx\mathbf{A}^{-1}$ is the preconditioner, $\mathbf{u}$ is the new unknown, and $\mathbf{x}=\mathbf{M}\mathbf{u}$ is the recovered solution. We design a graph neural network (GNN) [35, 33] to assume the role of $\mathbf{M}$ and develop a training method to learn $\mathbf{M}$ from select $(\mathbf{b},\mathbf{x})$ pairs. This proposal is inspired by the universal approximation property of neural networks [12] and encouraged by their widespread success in artificial intelligence [2]. We use a GNN to approximate the mapping from the right-hand side $\mathbf{b}$ to the solution $\mathbf{x}$.
@@
Because a neural network is, in nature, nonlinear, our preconditioner $\mathbf{M}$ is not a linear operator anymore. We can no longer build a Krylov subspace for $\mathbf{A}\mathbf{M}$ when a Krylov subspace is defined for only linear operators. Hence, we focus on flexible variants of the preconditioned Krylov methods instead. Specifically, we use flexible GMRES (FGMRES) [29], which considers $\mathbf{M}$ to be different in every Arnoldi step. In this framework, all what matters is the subspace from which the approximate solution is extracted. A nonlinear operator $\mathbf{M}$ can also be used to build this subspace and hence the neural preconditioner is applicable to FGMRES.
@@ 1.1 Contributions
Second, we propose an effective approach to training the neural preconditioner; in particular, training data generation. While it is straightforward to train the neural network in an online-learning manner—through preparing streaming, random input-output pairs in the form of $(\mathbf{b},\mathbf{x})$—it leaves open the question regarding what randomness best suits the purpose. We consider the bottom eigen-subspace of $\mathbf{A}$ when defining the sampling distribution.
@@
Fourth, we adopt a novel evaluation protocol for general-purpose preconditioners. The common practice evaluates a preconditioner on a limited number of problems; in many occasions, these problems come from a certain type of PDEs or an application, which constitutes only a portion of the use cases of solving linear systems. To broadly evaluate a new kind of general-purpose preconditioner and understand its strengths and weaknesses, we propose to test it on a large number of matrices commonly used by the community. To this end, we use the SuiteSparse collection and perform evaluation on a substantial portion of it (all non-SPD, real matrices falling inside a size interval; over 800 in this work). To streamline the evaluation, we define evaluation criteria, limit the tuning of hyperparameters, and study statistics over the distribution of matrices.
@@ 1.2 Related Work
It is important to put the proposed GNP in context. The idea of using a neural network to construct a preconditioner emerged recently. The relevant approaches learn the nonzero entries of the incomplete LU/Cholesky factors [20, 11] or of the approximate inverse [3]. In these approaches, the neural network as a function does not directly approximate the mapping from $\mathbf{b}$ to $\mathbf{x}$ like ours does; rather, it is used to complete the nonzeros of the predefined sparsity structure.
Approaches wherein the neural network directly approximates the matrix inverse are more akin to physics-informed neural networks (PINNs) [26] and neural operators (NOs) [17, 21, 22, 23]. However, the learning of PINNs requires a PDE (i.e., the physics) whose residual forms a training loss, whereas NOs consider infinite-dimensional function spaces, with PDEs being the primary application. In our case of a general-purpose preconditioner, there is not an underlying PDE; and not every matrix problem relates to function spaces with additional properties to exploit (e.g., spatial coordinates, smoothness, and decay of the Green’s function). For example, one may be interested in finding the commute times between a node and all other nodes in a graph; this problem can be solved by solving a linear system with respect the graph Laplacian matrix. The only information we exploit is the matrix itself.

### v3

@@ Abstract
Preconditioning is at the heart of iterative solutions of large, sparse linear systems of equations in scientific disciplines. Several algebraic approaches, which access no information beyond the matrix itself, are widely studied and used, but ill-conditioned matrices remain very challenging. We take a machine learning approach and propose using graph neural networks as a general-purpose preconditioner. They show attractive performance for many problems and can be used when the mainstream preconditioners perform poorly. Empirical evaluation on over 800 matrices suggests that the construction time of these graph neural preconditioners (GNPs) is more predictable and can be much shorter than that of other widely used ones, such as ILU and AMG, while the execution time is faster than using a Krylov method as the preconditioner, such as in inner-outer GMRES. GNPs have a strong potential for solving large-scale, challenging algebraic problems arising from not only partial differential equations, but also economics, statistics, graph, and optimization, to name a few.
@@ 1 Introduction
Iterative methods are commonly used to solve large, sparse linear systems of equations $\mathbf{A}\mathbf{x}=\mathbf{b}$. These methods typically build a Krylov subspace, onto which the original system is projected, such that an approximate solution within the subspace is extracted. For example, GMRES (Saad & Schultz, 1986), one of the most popularly used methods in practice, builds an orthonormal basis $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\}$ of the $m$-dimensional Krylov subspace by using the Arnoldi process and defines the approximate solution $\mathbf{x}_{m}\in\mathbf{x}_{0}+\operatorname{span}(\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{m}\})$, such that the residual norm $\|\mathbf{b}-\mathbf{A}\mathbf{x}_{m}\|_{2}$ is minimized.
@@
The effectiveness of Krylov methods heavily depends on the conditioning of the matrix $\mathbf{A}$. Hence, designing a good preconditioner is crucial in practice. In some cases (e.g., solving partial differential equations, PDEs), the problem structure provides additional information that aids the development of a preconditioner applicable to this problem or a class of similar problems. For example, multigrid preconditioners are particularly effective for Poisson-like problems with a mesh. In other cases, little information is known beyond the matrix itself. An example is the SuiteSparse matrix collection (Davis & Hu, 2011), which contains thousands of sparse matrices for benchmarking numerical linear algebra algorithms. In these cases, a general-purpose (also called “algebraic”) preconditioner is desirable; examples include ILU (Saad, 1994), approximate inverse (Chow & Saad, 1998), and algebraic multigrid (AMG) (Ruge & Stüben, 1987). However, it remains very challenging to design one that performs well universally.
@@
In this work, we propose to use neural networks as a general-purpose preconditioner. Specifically, we consider preconditioned Krylov methods for solving problems in the form $\mathbf{A}\mathbf{M}\mathbf{u}=\mathbf{b}$, where $\mathbf{M}\approx\mathbf{A}^{-1}$ is the preconditioner, $\mathbf{u}$ is the new unknown, and $\mathbf{x}=\mathbf{M}\mathbf{u}$ is the recovered solution. We design a graph neural network (GNN) (Zhou et al., 2020; Wu et al., 2021) to assume the role of $\mathbf{M}$ and develop a training method to learn $\mathbf{M}$ from select $(\mathbf{b},\mathbf{x})$ pairs. This proposal is inspired by the universal approximation property of neural networks (Hornik et al., 1989) and encouraged by their widespread success in artificial intelligence (Bommasani et al., 2021).
@@
Because a neural network is, in nature, nonlinear, our preconditioner $\mathbf{M}$ is not a linear operator anymore. We can no longer build a Krylov subspace for $\mathbf{A}\mathbf{M}$ when a Krylov subspace is defined for only linear operators. Hence, we focus on flexible variants of the preconditioned Krylov methods instead. Specifically, we use flexible GMRES (FGMRES) (Saad, 1993), which considers $\mathbf{M}$ to be different in every Arnoldi step. In this framework, all what matters is the subspace from which the approximate solution is extracted. A nonlinear operator $\mathbf{M}$ can also be used to build this subspace and hence the neural preconditioner is applicable to FGMRES.
We use a graph neural network because of the natural connection between a sparse matrix and the graph adjacency matrix, similar to how AMG interprets the coefficient matrix with a graph. Many GNN architectures access the graph structure through $\mathbf{A}$-multiplications and they become polynomials of $\mathbf{A}$ when the nonlinearity is omitted (Wu et al., 2019; Chen et al., 2020). This observation bridges the connection between a GNN as an operator and the polynomial approximation of $\mathbf{A}^{-1}$, the latter of which is an essential ingredient in the convergence theory of Krylov methods. Interestingly, a side benefit resulting from the access pattern of $\mathbf{A}$ is that the GNN is applicable to not only sparse matrices, but also structured matrices that admit fast $\mathbf{A}$-multiplications (such as hierarchical matrices (Hackbusch, 1999; Hackbusch & Börm, 2002; Chen & Stein, 2023), Toeplitz matrices (Chan & Jin, 2007; Chen et al., 2014), unassembled finite-element matrices (Ciarlet, 2002), and Gauss-Newton matrices (Liu et al., 2023)).
@@ 1.1 Contributions
Second, we propose an effective approach to training the neural preconditioner; in particular, training data generation. While it is straightforward to train the neural network in an online-learning manner—through preparing streaming, random input-output pairs in the form of $(\mathbf{b},\mathbf{x})$—it leaves open the question regarding what randomness best suits the purpose. We consider the bottom eigen-subspace of $\mathbf{A}$ when defining the sampling distribution and show the effectiveness of this approach.
@@
Fourth, we adopt a novel evaluation protocol for general-purpose preconditioners. The common practice evaluates a preconditioner on a limited number of problems; in many occasions, these problems come from a certain type of PDEs or an application, which constitutes only a portion of the use cases. To broadly evaluate a new kind of general-purpose preconditioner and understand its strengths and weaknesses, we propose to test it on a large number of matrices commonly used by the community. To this end, we use the SuiteSparse collection and perform evaluation on a substantial portion of it (all non-SPD, real matrices falling inside a size interval; over 800 from 50 application areas in this work). To streamline the evaluation, we define evaluation criteria, limit the tuning of hyperparameters, and study statistics over the distribution of matrices.
@@ 1.2 Related Work
It is important to put the proposed GNP in context. The idea of using a neural network to construct a preconditioner emerged recently. The relevant approaches learn the nonzero entries of the incomplete factors (Li et al., 2023; Häusner et al., 2023) or their correction (Trifonov et al., 2024), or of the approximate inverse (Bånkestad et al., 2024). In these approaches, the neural network does not directly approximate the mapping from $\mathbf{b}$ to $\mathbf{x}$ like ours does; rather, it is used to complete the nonzeros of the predefined sparsity structure. A drawback of these approaches is that the predefined sparsity imposes a nonzero lower bound on the approximation of $\mathbf{A}^{-1}$, which (and whose factors) are generally not sparse. Moreover, although the neural network can be trained on a PDE problem (by varying the coefficients, grid sizes, or boundary conditions) and generalizes well to the same problem, it is unlikely that a single network can work well for all problems and matrices in life.
Approaches wherein the neural network directly approximates the matrix inverse are more akin to physics-informed neural networks (PINNs) (Raissi et al., 2019) and neural operators (NOs) (Rudikov et al., 2024; Kovachki et al., 2022; Li et al., 2020; 2021; Lu et al., 2021). However, the learning of PINNs requires a PDE whose residual forms a partial training loss, whereas NOs consider infinite-dimensional function spaces, with PDEs being the application. In our case of a general-purpose preconditioner, there is not an underlying PDE; and not every matrix problem relates to function spaces with additional properties to exploit (e.g., spatial coordinates, smoothness, and decay of the Green’s function). For example, one may be interested in finding the commute times between a node and all other nodes in a graph; this problem can be solved by solving a linear system with respect the graph Laplacian matrix. The only information we exploit is the matrix itself.